OpenStudy (anonymous):

find the integral of: (3+4x)e^(6x+4x^2)

OpenStudy (lgbasallote):

\(\Large \int (3+4x) e^{(6x + 4x^2)} dx\) is that the question?

OpenStudy (anonymous):

yes, that is correct

OpenStudy (lgbasallote):

have you tried letting u = 6x + 4x^2?

OpenStudy (lgbasallote):

try it :) you'll find the answer :DDD i'll guide you

OpenStudy (anonymous):

Hey lgba u r great at it :D Congrats

OpenStudy (anonymous):

answer is 1/2 e^(2x(2x+3)) + c

OpenStudy (lgbasallote):

@oj4747 we do not give answers here at openstudy..we try to guide our askers as much as we can

OpenStudy (anonymous):

well that's what I thought but it was a complete guess...where do you get the 1/2?

OpenStudy (anonymous):

hmmm that from when u did subsitution

OpenStudy (lgbasallote):

like i said let u = 4x^2 + 3x du = (8x + 6)dx notice how if you divide both sides by two you get.. du/2 = (8x+6)/2 dx simplifying... du/2 = (4x + 3)dx notice that? so your integral would be \(\LARGE \int e^u \frac{du}{2}\) pull out 1/2 \(\LARGE \frac{1}{2} \int e^u du\) make sense?

OpenStudy (anonymous):

makes perfect sense...thanks!